OpenStudy (anonymous):

MEDAL REWARDED Find the inverse of the function. cuberoot x/6 -7

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

As we have, \[y=\sqrt[3]{x/6-7}\]Change x and y, \[x=\sqrt[3]{y/6-7}\]Then solve for y, \[x^3={y/6-7}\Rightarrow y=6x^3+42\]And that is the inverse.

OpenStudy (anonymous):

These are my answer choices: f-1(x) = 6(x+7)3 f-1(x) = 6(x3 + 7) f-1(x) = 18(x + 7) f-1(x) = [6(x+7)]3

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

Then it's the second, \[f^{-1}(x)=6x^3+42=6(x^3+7)\]

OpenStudy (anonymous):

Thank you so much

OpenStudy (john_es):

You're welcome ;).